OpenStudy (anonymous):

A who travels 3 miles an hour, starts from a certain place five hours in advance of B who travels 4 miles an hour in the same direction. How many hours must B travel to overtake A?

OpenStudy (radar):

When the distance traveled by A and B are equal then they have met and B then overtakes. Use d=rt (distance = rate times time.

OpenStudy (anonymous):

3t = 4(t-5) solve

OpenStudy (radar):

Note that 3t is distance of A and 4(t-5) is the distance of B.

OpenStudy (anonymous):

So how do I solve it?

OpenStudy (radar):

Are you an algebra student?

OpenStudy (anonymous):

yes, why? I know that distance = rate/time

OpenStudy (radar):

Use algebra and isolate t. Use algebraic rules and get t on the left and its' value on the right: t = ?? Hint clear the parenthesis of the equation provided by @sourwing

OpenStudy (anonymous):

3t = 4t-5

OpenStudy (radar):

and it is distance equals rate times time not divided by time.

OpenStudy (radar):

Now subtract 4t from both sides of the equal sign. also correct your error when you cleared the parenthesis.

OpenStudy (radar):

You should of gotten 3t = 4t - 20

OpenStudy (radar):

3t - 4t = 4t -4t -20

OpenStudy (radar):

-t = -20 Now divide both sides by -1

OpenStudy (anonymous):

Where did you get 20 from?

OpenStudy (anonymous):

Never mind

OpenStudy (anonymous):

It will take 20 hours to overtake A.

OpenStudy (radar):

Final results t = 20 hours. Should verify. Verification A distance 3 times 20 or 60 miles. B distnace 4(20-5) = 60 miles Yup, both traveled 60 miles.........good answer.

OpenStudy (radar):

Yes, cassieblake, 20 hours is good. And good luck with your studies.

OpenStudy (anonymous):

Thanks, Do you think you can help me with another problem?

OpenStudy (radar):

I can try.

OpenStudy (anonymous):

Say the rise is 4 and the run is 2. If I were to increase both the rise and run by 1, would it be the same rate of change? Why or why not?

OpenStudy (radar):

No, the rate of change would not remain the same. Here is why the original rise over run (rise/run) is 4/2 or 2. Now if we add 1 to both the rise (getting5) and the run (getting 3) or rise over run is 5/3. 5/3 does not = 2 the resulting change is at a lower rate.

OpenStudy (radar):

You follow ??

OpenStudy (anonymous):

okay, thanks.

OpenStudy (radar):

You're welcome.